The school drama club hosted a spring musical. The club charged $3 per child and $5 per adult. A total of 374 tickets were sold, raising 1,524. How many child and adult tickets were sold?( Please explain/ show work as well )
Answer by ikleyn(52810)

.
The school drama club hosted a spring musical. The club charged $3 per child and $5 per adult.
A total of 374 tickets were sold, raising 1,524. How many child and adult tickets were sold?
( Please explain/ show work as well )
~~~~~~~~~~~~~~~

Let x be the number of the adult tickets sold;

then the number of the child tickets is (374-x).


Write the total money equation (also called as revenue equation)

    5x + 3*(374-x) = 1524  dollars.


Simplify and find x

    5x + 3*374 - 3x = 1524

    5x - 3x = 1524 - 3*374

       2x   =     402

        x   =     402/2 

        x   =     201.


ANSWER.  201 adult tickets, and the rest, 374 - 201 = 173 are child tickets.


CHECK.   5*201 + 3*173 = 1524 dollars, in total.   ! Correct !

Solved.

-----------------

It is a standard and typical tickets problem.

There are different methods of solving such problems.
In this site, there is a lesson
    - Using systems of equations to solve problems on tickets
explaining and showing different methods of solving such problems.

From this lesson,  learn on how to solve such problems once and for all.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lesson is the part of this online textbook under the topic "Systems of two linear equations in two unknowns".

Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.